# Introducing Routing Functions to Vision-Language Parameter-Efficient Fine-Tuning with Low-Rank Bottlenecks

### v1

@@ Abstract
Mainstream parameter-efficient fine-tuning (PEFT) methods, such as LoRA or Adapter, project a model’s hidden states to a lower dimension, allowing pre-trained models to adapt to new data through this low-rank bottleneck. However, PEFT tasks involving multiple modalities, like vision-language (VL) tasks, require not only adaptation to new data but also learning the relationship between different modalities. Targeting at VL PEFT tasks, we propose a family of operations, called routing functions, to enhance VL alignment in the low-rank bottlenecks. The routing functions adopt linear operations and do not introduce new trainable parameters. In-depth analyses are conducted to study their behavior. In various VL PEFT settings, the routing functions significantly improve performance of the original PEFT methods, achieving over 20% improvement on VQAv2 ($\text{RoBERTa}_{\text{large}}$+ViT-L/16) and 30% on COCO Captioning (GPT2-medium+ViT-L/16). Also when fine-tuning a pre-trained multimodal model such as CLIP-BART, we observe smaller but consistent improvements across a range of VL PEFT tasks.
@@ 1 Introduction
When fine-tuning pre-trained foundation models for downstream tasks, studies show that a relatively high accuracy can be achieved by acting on a much smaller dimension compared to the original dimension of the pre-trained model’s feature space. The minimum dimension necessary to approximate the optimization problem of the downstream task is termed the intrinsic dimension li2018measuring ; aghajanyan-etal-2021-intrinsic .
@@
The existence of such intrinsic dimension explains the success of parameter-efficient fine-tuning (PEFT) methods with low-rank bottlenecks for model fine-tuning. Two of the most widely applied techniques are Adapter pmlr-v97-houlsby19a and LoRA hu2022lora . We demonstrate the use of Adapter and LoRA within a Transformer block in Footnote 2. Despite being added at different positions within the Transformer block, Adapter and LoRA exhibit some key characteristics in their architectural design. Specifically, they first map the hidden states $x_{H}$ to a lower dimension $r$ through a down-projection mapping $W_{down}$. The down-projected states are then projected back to the original dimension $d$ via an up-projection mapping $W_{up}$. In the PEFT context, only the parameters of the colored modules in Footnote 2 are updated, while the remaining parameters are unchanged. Since the intermediate dimension $r$ is much smaller than the original dimension $d$, these methods significantly reduce the number of parameters that need to be updated.
@@
Figure 1: Illustration of LoRA and Adapter used in a Transformer block. Only the green modules are updated in PEFT. We identify the low-rank bottleneck with the orange rectangle. The Adapters are sequentially added to the Transformer block.222 In this visualization, we omit the nonlinear function in the Adapter for simplicity.
@@
In uni-modal tasks, where $x_{H}$ contains information from a single modality, $W_{down}$ effectively compresses $x_{H}$ into a feature space of lower dimension. This compression usually suffices provided that $r$ is larger than the intrinsic dimension of the downstream task hu2022lora . However, PEFT methods with low-rank bottlenecks, such as Adapter and LoRA, are currently not designed for down-stream tasks involving another modality, such as vision-language (VL) tasks. Even for a pre-trained multimodal foundation model with $x_{H}$ containing information from both modalities, conceptually, directly compressing $x_{H}$ with $W_{down}$ poses challenges, as it is unclear how to effectively balance the two modalities and to enforce alignment between the modalities with this simple linear mapping. Therefore, in this paper we investigate whether it is possible to better route the features through the low-rank bottleneck with some extra guidance. An intuitive approach is standard cross-attention. However, it introduces additional parameters through query/key/value projections. Moreover, softmax-based operations not only increase time complexity but are limited to dealing with features of length $>$ 1. This is incompatible with the PEFT setting, where the goal is to minimize the number of trainable parameters and computational complexity, while maintaining versatility across different types of tasks (including those that use only the [CLS] features). In this work, we introduce routing functions in VL PEFT with low-rank bottlenecks. They rely on linear operations and do not introduce extra trainable parameters.
@@ 2.1 Parameter-Efficient Fine-Tuning (PEFT)
Adapter-based PEFT To reduce the tuning costs of pre-trained language models (LMs), Adapter was first introduced in pmlr-v97-houlsby19a . The Adapter layer first down-projects the model’s hidden state, then up-projects it back to its original dimension, incorporating a nonlinear function after the down-projection. Freezing the pre-trained weights of the LM and tuning them with sequentially added Adapter layers achieves performance comparable to full fine-tuning. Since then, various Adapter variants have been proposed to advance the field of PEFT. The focus of developing Adapter variants has been on improving efficiency ruckle-etal-2021-Adapterdrop ; mahabadi2021compacter ; lei2023conditional and exploring alternative architectures lin-etal-2020-exploring ; karimi-mahabadi-etal-2021-parameter ; he2022towards . Despite being surpassed by its variants on various benchmarks, the original sequential Adapter is still widely used in practice and offers stable results across tasks.
@@
LoRA-based PEFT LoRA hu2022lora adopts the same "down-projection then up-projection" approach as the Adapter, with no nonlinear function in the low-rank bottleneck. LoRA is added to the query and value mappings in the attention modules of the pre-trained LM. Important follow-up works include dynamic rank adjustment valipour-etal-2023-dylora ; zhang2023adaptive and quantization adaptation NEURIPS2023_qlora ; xu2024qalora . Moreover, VeRA kopiczko2024vera and NOLA koohpayegani2024nola further minimize the number of parameters by leveraging frozen random matrices. LoRA-based PEFT methods have become standard for tuning large language models and large multimodal foundation models.
@@
Others Beyond Adapter and LoRA, prompt-based tuning methods lester-etal-2021-power ; li-liang-2021-prefix ; jia2022vpt have shown strong performance for tuning both language and vision models. LST sung2022lst learns a compact side network during tuning. BitFit ben-zaken-etal-2022-bitfit only adjusts the bias terms during fine-tuning, and achieves comparable performance as other PEFT methods. Methods mao-etal-2022-unipelt ; zhou-etal-2024-autopeft ; he2022towards ; wang-etal-2022-adamix also explore combining PEFT methods to achieve superior performance.
2.2 Vision-Language PEFT
@@
VL-Adapter Sung_2022_CVPR is the pioneering work in the field of Vision-Language (VL) PEFT. It empirically studies the use of different PEFT methods on four different VL tasks in a multitask learning setting. VL-Adapter uses CLIP-BART, a similar model architecture as VL-T5 2020t5 . CLIP-BART uses a projection layer to connect CLIP pmlr-v139-radford21a vision encoder to BART lewis-etal-2020-bart or T5 2020t5 . CLIP-BART is tuned using a combination of prompt tuning and other PEFT methods. We conduct extensive experiments using CLIP-BART and VL-Adapter as presented in Sec. 5. As an extension to VL-Adapter, VL-PET Hu_2023_ICCV combines an Adapter with a rather complicated granularity-controlled mechanism. I-Tuning ituning2023 explores alternative model architectures for VL PEFT tasks. and proposes a parallel Adapter with cross-attention to integrate visual information into a pre-trained LM. Since our work is the first that introduces routing functions into VL PEFT, we limit the comparison of our method to VL-Adapter and I-Tuning, leaving the exploration of more complicated PEFT methods like VL-PET to future work.
@@ 3 Methodology
As illustrated in Figure 2 (Left), standard PEFT modules with low-rank bottleneck, i.e. LoRA/Adapter, learn a down-projection mapping $W_{down}$ followed by an up-projection mapping $W_{up}$, with no routing in between. These modules are added in each Transformer block of the pre-trained foundation model (Footnote 2). As presented in Figure 2 (Middle), our method adds a routing function in the low-rank bottleneck. This routing function is designed to align two sets of features, in particular, the extra visual features $x_{R}$ to be integrated in the VL tasks and the hidden states $x_{H}$ of each Transformer block of the foundation model.
@@
In this paper, we explore four types of routing functions. Given $x_{H}$ and $x_{R}$ [^3] Assuming $x_{R}$ to have the same dimension $d$ as $x_{H}$. See Sec. 4Sec. 5 for details., after the $W_{down}$ mapping, we obtain $x_{t}=W_{down}x_{H}$ and $x_{v}=W_{down}x_{R}$. Here $x_{t}$ is of shape $(L_{t},r)$ and $x_{v}$ is of shape $(L_{v},r)$, where $L_{t}$/$L_{v}$ and $r$ are the feature length (i.e. the number of tokens) and the bottleneck dimension or rank, respectively. [^4] Here we omit the batch dimension for simplicity. In practice, we always use $L_{v}=1$, i.e. we use a single token / global feature vector to represent the visual content. In case of multiple tokens, we use average pooling to reduce it to a single feature vector. [^5] Only when comparing to cross-attention in Sec. 4.3, we do not use pooled features. The proposed routing functions operate as:
@@ 4.1 Experimental Design
Here, we study VL PEFT tasks in their cleanest form. Specifically, (1) No additional mapping networks are adopted, so the routing functions form the only module that contributes to the final performance; (2) Only the parameters of the PFET modules in LMs are updated, where we learn the VL alignment; (3) No multimodal tasks (e.g. contrastive language-image pre-training in CLIP pmlr-v139-radford21a ) are included in the pre-training phases of the backbone models. We only learn the multimodal embedding space during VL PEFT. These principles guarantee that VL alignment originates strictly from the PEFT modules.
@@
To comprehensively study the role of routing functions in VL tasks, we consider both discriminative and generative VL tasks, to comprehensively study the role of different routing functions. We select visual question answering (VQA) and Image Captioning, two of the most widely studied VL tasks, as the representative task for discriminative VL tasks and generative VL tasks, respectively.
We present the pipelines for VQA and Image Captioning in Figure 3. We use ViT dosovitskiy2021an as our vision backbone. As for the LM backbone, we use the encoder-only RoBERTa roberta19 and the decoder-only GPT2 radford2019language for VQA and Image Captioning, respectively. [^7] In Appendix 0.B we report results obtained with larger backbones, e.g. $\text{RoBERTa}_{\text{large}}$. The [CLS] feature from the ViT is prepended to the text inputs and acts as $x_{R}$, enabling the learning of VL alignment through PEFT, with or without the integration of our routing functions. Prepending visual features to text inputs is actually the predominant way of learning VL alignment in PEFT Sung_2022_CVPR ; Hu_2023_ICCV without routing functions. The ViT and LMs adopt the same dimensions for their feature spaces, so we do not need additional mappings to match the dimensions.
@@ 4.2 Implementation Details
We adopt ViT-B/16 pre-trained on ImageNet-21k as our vision backbone, which we keep frozen during training. In LoRA, we set scaling factor $\alpha=r$ in all setups. We use the Karpathy split of VQAv2 balanced_vqa_v2 and COCO Captioning mscoco (COCO Cap.) in our experiments. For VQAv2, we use $\text{RoBERTa}_{\text{base}}$ as our LM backbone, with a two-layer multi-layer perceptron (MLP) classifier on top following Sung_2022_CVPR . Only the parameters of the classifier and the PEFT modules in RoBERTa are updated during training. We train our model using the AdamW optimizer loshchilov2018decoupled with learning rate=3e-4 and batch size of 512 for 20 epochs, where we warm up the first 5% of the steps. For the captioning task, COCO Cap., we use GPT2-base as our LM backbone. Only the parameters of the PEFT modules in GPT2 are updated during training. The maximum length for captions is set to 50. We train our model for 20 epochs using the AdamW optimizer with batch size of 40 and learning rate=3e-5, where we warm up the first 5000 steps. After training, we evaluate the models using the final checkpoint for both tasks.
@@ 4.3 Results
$x_{t}\circ x^{\prime}_{v}$ $9.3653.51_{\color[rgb]{0.1914,0.3359375,0.9414}{+9.36}}$ $8.6252.78_{\color[rgb]{0.1914,0.3359375,0.9414}{+8.62}}$ $10.41\mathbf{53.86_{\color[rgb]{0.1914,0.3359375,0.9414}{+10.41}}}$ $8.7353.01_{\color[rgb]{0.1914,0.3359375,0.9414}{+8.73}}$
$x_{t}+x^{\prime}_{v}$ $8.4552.60_{\color[rgb]{0.1914,0.3359375,0.9414}{+8.45}}$ $9.7853.94_{\color[rgb]{0.1914,0.3359375,0.9414}{+9.78}}$ $8.4352.88_{\color[rgb]{0.1914,0.3359375,0.9414}{+8.43}}$ $9.6753.95_{\color[rgb]{0.1914,0.3359375,0.9414}{+9.67}}$
$x_{t}(x_{v})^{\mathsf{T}}x_{v}$ $9.73\underline{53.88_{\color[rgb]{0.1914,0.3359375,0.9414}{+9.73}}}$ $10.32\underline{54.48_{\color[rgb]{0.1914,0.3359375,0.9414}{+10.32}}}$ $8.64\underline{53.09_{\color[rgb]{0.1914,0.3359375,0.9414}{+8.64}}}$ $10.78\mathbf{55.06_{\color[rgb]{0.1914,0.3359375,0.9414}{+10.78}}}$
$x_{t}x^{\prime\prime}_{v}$ $10.06\mathbf{54.21_{\color[rgb]{0.1914,0.3359375,0.9414}{+10.06}}}$ $10.80\mathbf{54.96_{\color[rgb]{0.1914,0.3359375,0.9414}{+10.80}}}$ $7.4351.88_{\color[rgb]{0.1914,0.3359375,0.9414}{+7.43}}$ $\underline{\color[rgb]{0.1914,0.3359375,0.9414}{54.38_{+10.10}}}$
Without the addition of routing functions, models with LoRA and Adapter reach a similar level of performance. All four types of routing functions significantly improve the performance of the original PEFT method in terms of accuracy of the answer. The dimension of the bottleneck does not have a strong impact on the final performance. Sometimes we can even obtain better performance with lower dimensions. It is a sign that the intrinsic dimension of the VL task is not high, just as observed in uni-modal tasks. Moreover, we observe that the routing functions involving matrix multiplications (i.e. $x_{t}(x_{v})^{\mathsf{T}}x_{v}$ and $x_{t}x^{\prime\prime}_{v}$) perform quite well in most scenarios for both LoRA and Adapter, while the routing functions involving element-wise operations (i.e. $x_{t}\circ x^{\prime}_{v}$ and $x_{t}+x^{\prime}_{v}$) are just slightly behind.

### v2

@@ Abstract
Mainstream parameter-efficient fine-tuning (PEFT) methods, such as LoRA or Adapter, project a model’s hidden states to a lower dimension, allowing pre-trained models to adapt to new data through this low-rank bottleneck. However, PEFT tasks involving multiple modalities, like vision-language (VL) tasks, require not only adaptation to new data but also learning the relationship between different modalities. Targeting at VL PEFT tasks, we propose a family of operations, called routing functions, to enhance VL alignment in the low-rank bottlenecks. These feature routing functions adopt linear operations and do not introduce new trainable parameters. In-depth analyses are conducted to study their behavior. In various VL PEFT settings, the routing functions significantly improve performance of the original PEFT methods, achieving over 20% improvement on VQAv2 ($\text{RoBERTa}_{\text{large}}$+ViT-L/16) and 30% on COCO Captioning (GPT2-medium+ViT-L/16). Also when fine-tuning a pre-trained multimodal model such as CLIP-BART, we observe smaller but consistent improvements across a range of VL PEFT tasks. Our code is available at https://github.com/tingyu215/Routing_VLPEFT.
@@ 1 Introduction
When fine-tuning pre-trained foundation models for downstream tasks, studies show that a relatively high accuracy can be achieved by acting on a much smaller dimension compared to the original dimension of the pre-trained model’s feature space. The minimum dimension necessary to approximate the optimization problem of the downstream task is termed the intrinsic dimension [21, 1].
@@
The existence of such intrinsic dimension explains the success of parameter-efficient fine-tuning (PEFT) methods with low-rank bottlenecks for model fine-tuning. Two of the most widely applied techniques are Adapter [9] and LoRA [10]. We demonstrate the use of Adapter and LoRA within a Transformer block in Footnote 2. Despite being added at different positions within the Transformer block, Adapter and LoRA exhibit some key characteristics in their architectural design. Specifically, they first map the hidden states $x_{H}$ to a lower dimension $r$ through a down-projection mapping $W_{down}$. The down-projected states are then projected back to the original dimension $d$ via an up-projection mapping $W_{up}$. In the PEFT context, only the parameters of the colored modules in Footnote 2 are updated, while the remaining parameters are unchanged. Since the intermediate dimension $r$ is much smaller than the original dimension $d$, these methods significantly reduce the number of parameters that need to be updated.
@@
Figure 1: Illustration of LoRA and Adapter used in a Transformer block. Only the green modules are updated in PEFT. We identify the low-rank bottleneck with the orange rectangle. The Adapters are sequentially added to the Transformer block.22footnotemark: 2
@@
In uni-modal tasks, where $x_{H}$ contains information from a single modality, $W_{down}$ effectively compresses $x_{H}$ into a feature space of lower dimension. This compression usually suffices provided that $r$ is larger than the intrinsic dimension of the downstream task [10]. However, PEFT methods with low-rank bottlenecks, such as Adapter and LoRA, are currently not designed for down-stream tasks involving another modality, such as vision-language (VL) tasks. Even for a pre-trained multimodal foundation model with $x_{H}$ containing information from both modalities, conceptually, directly compressing $x_{H}$ with $W_{down}$ poses challenges, as it is unclear how to effectively balance the two modalities and to enforce alignment between the modalities with this simple linear mapping. Therefore, in this paper we investigate whether it is possible to better route the features through the low-rank bottleneck with some extra guidance. An intuitive approach is standard cross-attention. However, it introduces additional parameters through query/key/value projections. Moreover, softmax-based operations not only increase time complexity but are limited to dealing with features of length $>$ 1. This is incompatible with the PEFT setting, where the goal is to minimize the number of trainable parameters and computational complexity, while maintaining versatility across different types of tasks (including those that use only the [CLS] features). In this work, we introduce feature routing functions in VL PEFT with low-rank bottlenecks. They rely on linear operations and do not introduce extra trainable parameters.
@@ 2 Related Works
Adapter-based PEFT To reduce the tuning costs of pre-trained language models (LMs), the Adapter layer was first introduced in [9]. It first down-projects the model’s hidden state, then up-projects it back to its original dimension, incorporating a nonlinear function after the down-projection. Freezing the pre-trained weights of the LM and tuning them with sequentially added Adapter layers achieves performance comparable to full fine-tuning. Since then, various Adapter variants have been proposed to advance the field of PEFT. The focus of developing Adapter variants has been on improving efficiency [34, 28, 18] and exploring alternative architectures [24, 14, 8]. Despite being surpassed by its variants on various benchmarks, the original sequential Adapter is still widely used in practice and offers stable results across tasks.
@@
LoRA-based PEFT LoRA [10] adopts the same "down-projection then up-projection" approach as the Adapter, with no nonlinear function in the low-rank bottleneck. LoRA is added to the query and value mappings in the attention modules of the pre-trained LM. Important follow-up works include dynamic rank adjustment [38, 41] and quantization adaptation [5, 40]. Moreover, VeRA [17] and NOLA [16] further minimize the number of parameters by leveraging frozen random matrices. LoRA-based PEFT methods have become standard for tuning large language models and large multimodal foundation models.
@@
Other PEFT methods Beyond Adapter and LoRA, prompt-based tuning methods [19, 22, 13] have shown strong performance for tuning both language and vision models. LST [36] learns a compact side network during tuning. BitFit [2] only adjusts the bias terms during fine-tuning, and achieves comparable performance as other PEFT methods. Methods [29, 42, 8, 39] also explore combining PEFT methods to achieve superior performance.
@@
Vision-Language PEFT VL-Adapter [37] is the pioneering work in the field of Vision-Language (VL) PEFT. It empirically studies the use of different PEFT methods on four different VL tasks in a multitask learning setting. VL-Adapter uses CLIP-BART, a similar model architecture as VL-T5 [33]. CLIP-BART uses a projection layer to a connect CLIP [31] vision encoder to BART [20] or T5 [33]. CLIP-BART is tuned using a combination of prompt tuning and other PEFT methods. We conduct extensive experiments using CLIP-BART and VL-Adapter as presented in Sec. 5. As an extension to VL-Adapter, VL-PET [11] combines an Adapter with a rather complicated granularity-controlled mechanism. I-Tuning [27] explores alternative model architectures for VL PEFT tasks. and proposes a parallel Adapter with cross-attention to integrate visual information into a pre-trained LM. Since our work is the first that introduces routing functions into VL PEFT, we limit the comparison of our method to VL-Adapter and I-Tuning, leaving the exploration of more complicated PEFT methods like VL-PET to future work.
@@ 3 Methodology
As illustrated in Fig. 2 (Left), standard PEFT modules with low-rank bottleneck, i.e. LoRA/Adapter, learn a down-projection mapping $W_{down}$ followed by an up-projection mapping $W_{up}$, with no routing in between. These modules are added in each Transformer block of the pre-trained foundation model (Footnote 2). As presented in Fig. 2 (Middle), our method adds a routing function in the low-rank bottleneck. This routing function is designed to align two sets of features, in particular, the extra visual features $x_{R}$ to be integrated in the VL tasks and the hidden states $x_{H}$ of each Transformer block of the foundation model.
@@
In this paper, we explore four types of routing functions. Given $x_{H}$ and $x_{R}$ [^3] Assuming $x_{R}$ to have the same dimension $d$ as $x_{H}$. See Sec. 4 and Sec. 5 for details., after the $W_{down}$ mapping, we obtain $x_{t}=W_{down}x_{H}$ and $x_{v}=W_{down}x_{R}$. Here $x_{t}$ is of shape $(L_{t},r)$ and $x_{v}$ is of shape $(L_{v},r)$, where $L_{t}$/$L_{v}$ and $r$ are the feature length (i.e. the number of tokens) and the bottleneck dimension or rank, respectively. [^4] Here we omit the batch dimension for simplicity. In practice, we always use $L_{v}=1$, i.e. we use a single token / global feature vector to represent the visual content. In case of multiple tokens, we use average pooling to reduce it to a single feature vector. [^5] Only when comparing to cross-attention in Sec. 4.3, we do not use pooled features. The proposed routing functions operate as:
@@ 4.1 Experimental Design
Here, we study VL PEFT tasks in their cleanest form. Specifically, (1) No additional mapping networks are adopted, so the feature routing functions form the only module that contributes to the final performance; (2) Only the parameters of the PFET modules in LMs are updated, where we learn the VL alignment; (3) No multimodal tasks (e.g. contrastive language-image pre-training in CLIP [31]) are included in the pre-training phases of the backbone models. We only learn the multimodal embedding space during VL PEFT. These principles guarantee that VL alignment originates strictly from the PEFT modules.
@@
To comprehensively study the role of routing functions in VL tasks, we consider both discriminative and generative VL tasks. We select visual question answering (VQA) and Image Captioning, two of the most widely studied VL tasks, as the representative task for discriminative VL tasks and generative VL tasks, respectively.
We present the pipelines for VQA and Image Captioning in Fig. 3. We use ViT [6] as our vision backbone. As for the LM backbone, we use the encoder-only RoBERTa [25] and the decoder-only GPT2 [32] for VQA and Image Captioning, respectively.777 In Appendix 0.B, we report results obtained with larger backbones, e.g. $\text{RoBERTa}_{\text{large}}$. The [CLS] feature from the ViT is prepended to the text inputs and acts as $x_{R}$, enabling the learning of VL alignment through PEFT, with or without the integration of our routing functions. Prepending visual features to text inputs is actually the predominant way of learning VL alignment in PEFT [37, 11] without routing functions. The ViT and LMs adopt the same dimensions for their feature spaces, so we do not need additional mappings to match the dimensions.
@@ 4.2 Implementation Details
We adopt ViT-B/16 pre-trained on ImageNet-21k as our vision backbone, which we keep frozen during training. In LoRA, we set scaling factor $\alpha=r$ in all setups. For VQAv2, we use $\text{RoBERTa}_{\text{base}}$ as our LM backbone, with a two-layer multi-layer perceptron (MLP) classifier on top following [37]. We train our model using the AdamW optimizer [26] with learning rate=3e-4 and batch size of 512 for 20 epochs, where we warm up the first 5% of the steps. For the captioning task, COCO Cap., we use GPT2-base as our LM backbone. We train our model for 20 epochs using the AdamW optimizer with batch size of 40 and learning rate=3e-5, where we warm up the first 5000 steps. For more details, see Appendix 0.A.
@@ 4.3 Results
$x_{t}\circ x^{\prime}_{v}$ $9.3653.51_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+9.36}}$ $8.6252.78_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+8.62}}$ $10.41\mathbf{53.86_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+10.41}}}$ $8.7353.01_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+8.73}}$
$x_{t}+x^{\prime}_{v}$ $8.4552.60_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+8.45}}$ $9.7853.94_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+9.78}}$ $8.4352.88_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+8.43}}$ $9.6753.95_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+9.67}}$
$x_{t}(x_{v})^{\mathsf{T}}x_{v}$ $9.73\underline{53.88_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+9.73}}}$ $10.32\underline{54.48_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+10.32}}}$ $8.64\underline{53.09_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+8.64}}}$ $10.78\mathbf{55.06_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+10.78}}}$
$x_{t}x^{\prime\prime}_{v}$ $10.06\mathbf{54.21_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+10.06}}}$ $10.80\mathbf{54.96_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+10.80}}}$ $7.4351.88_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+7.43}}$ $\underline{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{54.38_{+10.10}}}$
Without the addition of routing functions, models with LoRA and Adapter reach a similar level of performance. All four types of routing functions significantly improve the performance of the original PEFT method in terms of accuracy of the answer. The dimension of the bottleneck does not have a strong impact on the final performance. Sometimes we can even obtain better performance with lower dimensions. It is a sign that the intrinsic dimension of the VL task is not high, just as observed in uni-modal tasks. Moreover, we observe that the routing functions involving matrix multiplications (i.e. $x_{t}(x_{v})^{\mathsf{T}}x_{v}$ $x_{t}x^{\prime\prime}_{v}$) perform quite well in most scenarios for both LoRA and Adapter, while the routing functions involving element-wise operations (i.e. $x_{t}\circ x^{\prime}_{v}$ $x_{t}+x^{\prime}_{v}$) are just slightly behind.